OpenStudy (anonymous):

A line contains the points (3, -1) and (-1, 2). Another line graphed in the same coordinate plane contains the points (2,0) and (-2,3). Based on the slope of these lines, are they parallel, perpendicular or neither?

OpenStudy (amistre64):

what are their slopes?

OpenStudy (amistre64):

the question asks "based on their slopes" which implies that the slopes are already found. So what have you found their slopes to be?

OpenStudy (anonymous):

3/4 and 3/-4

OpenStudy (amistre64):

good, now ill dbl chk that to be sure :) (3, -1) -(-1, 2) -------- 4,-3 ; -3/4 (2,0) -(-2,3) ------- 4,-3 ; -3/4

OpenStudy (amistre64):

they have the exact same slope; when 2 lines have the same slope, that defines parallel lines.

OpenStudy (amistre64):

how did you find the slope? just so i can see what went awry for you

OpenStudy (anonymous):

i used the slope formula and i subtract 2-(-1) and i subt 3-(-1) and i found the answer for 3/4

OpenStudy (amistre64):

the formula is good, but it tends to get confusing trying to make sure that the parts are in the right spots. in essense; you subtract one point from another: (3, -1) -(-1, 2) (3, -1) 1, -2 -------- 4, -3 ; and stack y/x. slope = -3/4 and again for the other one: (2,0) -(-2,3) (2,0) 2,-3 ------- 4,-3; slope = y/x, -3/4

OpenStudy (amistre64):

i just find that "method" easier to follow

OpenStudy (anonymous):

it look easier thank you

OpenStudy (amistre64):

youre welcome, and good luck :)

OpenStudy (anonymous):

thnkz